If eight identical drops are joined to form a bigger drop, the potential on bigger as compared to that on smaller drop will be ?

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
Here we accept that r and R be the radii of the individual mercury drop and the mixed 
drop. 
Volume of the blended drop = 4πR3/3 = 8 x 4πr3/3 
⇒ R = 2r. 
Give q a chance to be the charge on every mercury drop and Q = 8q be the charge on the blended drop. 
V = kq/r = 20 V 
V' = kQ/R 
= k. 8q/(2r) 
= 4 kq/r 
= 4 x 20 V = 80 V.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)